OpenStudy (simplixity):

A penny is placed a distance r from the center of a record spinning at ω = 90π rad/min. The coefficient of static friction between the penny and the record is μs = 0.14 on the horizontal plane. What is the distance, r in meters?

OpenStudy (3mar):

Well, I am here.

OpenStudy (simplixity):

The equation to find r is \[(\mu*g)/\omega^2\] but I'm not getting the right answer for some reason

OpenStudy (3mar):

Where are you stuck?

OpenStudy (simplixity):

getting the right answer

OpenStudy (3mar):

\[\frac{ \mu*g }{ \omega^2 } =??\]

OpenStudy (welshfella):

You need to be careful with the units

OpenStudy (simplixity):

So this is how I did it.. \[\frac{ 0.14 *9.81}{ 90\pi^2}\]

OpenStudy (welshfella):

that denominator is wrong

OpenStudy (3mar):

Hint: 1. convert min to sec. 2. square the w>>=\(w^2=90^2*pi^2\)

OpenStudy (3mar):

but that without conversion to rad/sec

OpenStudy (welshfella):

9.81 is in m/s-2 and speed is in in minutes

OpenStudy (welshfella):

convert to rads/sec

OpenStudy (3mar):

@simplixity

OpenStudy (simplixity):

soo would I do this? \[\frac{ 90\pi }{ 1 rev }*\frac{ 1 \min }{ 60 \sec }\] to convert?

OpenStudy (simplixity):

I got 4.712

OpenStudy (3mar):

Great! Sub. with it it the formula you typed!

OpenStudy (simplixity):

is this how it should look? \[\frac{ 0.14*4.712 }{ 90^2\pi^2 }\]

OpenStudy (3mar):

I am here!

OpenStudy (3mar):

No, I don't think so. It would be\[\Large \frac{ \mu*g }{ \omega^2 } =\frac{ (0.14)(9.81) }{ (4.7124)^2 }=0.06185\] I ask you know, what does this value represent?

OpenStudy (3mar):

Is there any graph for that problem?

OpenStudy (simplixity):

thank you for your help

OpenStudy (3mar):

Thank you for the medal! Did you get what you were looking for?

OpenStudy (simplixity):

no problem and yes, I understand it now

OpenStudy (3mar):

Any more questions?

OpenStudy (3mar):

I am happy to hear that.

OpenStudy (simplixity):

nope that's it. thank you, I appreciate it

OpenStudy (3mar):

I will not be late for any help.